OpenStudy (anonymous):

Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 14x - y = -10 7x + 2y = 15

OpenStudy (texaschic101):

Look at your first equation....we want to get y by itself. 14x - y = -10 --- add y to both sides 14x = -10 + y -- add 10 to both sides 14x + 10 = y now sub 14x + 10 in for y in the other equation 7x + 2y = 15 7x + 2(14x + 10) = 15 --- distribute through the parenthesis 7x + 28x + 20 = 15 -- subtract 20 from both sides 7x + 28x = 15 - 20 -- combine like terms 35x = -5 -- divide both sides by 35 x = -5/35 x = - 1/7 now sub -1/7 in for x in either equation to find y 14x - y = -10 14(-1/7) - y = -10 -14/7 - y = -10 -2 - y = -10 -y = -10 + 2 -y = -8 y = 8 check... 7x + 2y = 15 7(-1/7) + 2(8) = 15 -1 + 16 = 15 15 = 15 (correct) x = -1/7 and y = 8 any questions ?

OpenStudy (amistre64):

14x - y = -10 7x + 2y = 15 let x=(n/14 -5/14) let y=(n+5) and substitute 14(n/14 - 5/14) - (n+5) = -10 7(n/14 - 5/14) + 2(n+5) = 15 n - 5 -n -5 = -10 n/2 - 5/2 + 2n +10 = 15 - 5-5 = -10 ; always true 5n/2 - 15/2 = 15; solve for n

OpenStudy (amistre64):

prolly typoed somethign in there :)

OpenStudy (anonymous):

@texaschic101 Thank You for explaining. It really helped

OpenStudy (texaschic101):

I am always glad to help :)